OpenStudy (anonymous):

NEED HELP ASAP!! Let {u1, u2,u3} be a linearly independent set of vectors in R^n, and let vector v be a vector in R^n such that v= c1 u1+c2us+c3u3 for some scalars c1, c2,c3 wiht c3 not equal to 0. prove that the set {v, u1, u2} is linearly independent.

OpenStudy (anonymous):

v= c1 u1+c2us+c3u3 is it c2 u2

OpenStudy (usukidoll):

argh I used to know this :*(

OpenStudy (anonymous):

V=C1U1 +C2 U2 + C3U3

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

Do you guys know how to do it?

OpenStudy (anonymous):

Yes, Stay tuned

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

Let @wio help you. My wife is telling me to go to dinner

OpenStudy (anonymous):

Proof by contradiction might be a good way to go. If {v, u1, u2} is not linearly dependent, then {u1, u2,u3} is not.

OpenStudy (anonymous):

If {v, u1, u2} is not linearly dependent, then we have constants \(k_1,k_2\) such that \[ v = k_1u_1+k_2u_2 \]

OpenStudy (anonymous):

We know that \[ v = c_1 u_1+c_2u_2+c_3u_3 \implies \frac{1}{c_3}\left( v - c_1 u_1 -c_2u_2 \right)= u_3 \]

OpenStudy (anonymous):

Now, substitute \(v=k_1u_1+k_2u_2\). Does it make sense?

OpenStudy (anonymous):

I suppose this is actually a proof by contrapositive.

OpenStudy (anonymous):

ok thank you :)

OpenStudy (anonymous):

So you can take it from here? You know where I'm going? Good!

OpenStudy (anonymous):

Notice this proof wouldn't work if we didn't know beforehand that \(c_3\neq 0\).

OpenStudy (anonymous):

@UsukiDoll I winged this one.

OpenStudy (usukidoll):

yay!

OpenStudy (anonymous):

When I made the first post, I didn't actually know what I was going to do, and I just sort of came up with it as I went along.

OpenStudy (usukidoll):

XD

OpenStudy (anonymous):

im sorry im actually stuck

OpenStudy (anonymous):

No problem. Where are you stuck?

OpenStudy (usukidoll):

copy word for word XD

OpenStudy (anonymous):

after substituting v= k1u1+K2u2

OpenStudy (anonymous):

Okay, well, here is a hint. We want to show that \(u_3\) is linearly dependent on \(u_1,u_2\). This means that there are constants such that \[ u_3=d_1u_1+d_2u_2 \]You just need to find \(d_1,d_2\).

OpenStudy (usukidoll):

\[v=k_1u_1+k_2u_2\]

OpenStudy (anonymous):

Do you understand how to find the \(d\)?

OpenStudy (anonymous):

no thats where im confused

OpenStudy (anonymous):

Okay so you have \[ u_3=\frac{1}{c_3}\left(k_1u_1+k_2u_2-c_1u_1-c_2u_2\right) \]

OpenStudy (anonymous):

combine like terms.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

(1/c3( k3u3-c1u1)

OpenStudy (anonymous):

idk if its correct

OpenStudy (anonymous):

Okay, so the problem is not just linear algebra, but also algebra.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\begin{split} u_3 &= \frac{1}{c_3}\left(k_1u_1+k_2u_2-c_1u_1-c_2u_2\right)\\ &= \frac{1}{c_3}\left((k_1-c_1)u_1+(k_2-c_2)u_2\right)\\ &=\left( \frac{k_1-c_1}{c_3}\right) u_1+\left(\frac{k_2-c_2}{c_3}\right)u_2\\ &=d_1 u_1+d_2u_2\\ \end{split}\]

OpenStudy (anonymous):

aa i completely forgot how to combine it. it looks more familiar with numbers. im sorry.

OpenStudy (anonymous):

now i have to find the values of d1 and d2

OpenStudy (anonymous):

No, this completes the proof. We have shown that they exists, which means that \(u_1,u_2,u_3\) are not linearly independent.

OpenStudy (anonymous):

ok I have to learn this topic more thoroughly. Doing an example helps. I have an exam tomorrow and I have no idea how to do this. Thank you for your help!! :D

OpenStudy (usukidoll):

I thought exams don't have proof problems D:

OpenStudy (usukidoll):

but the higher level courses do

OpenStudy (anonymous):

unfortuanly my teacher is asking for proofs

OpenStudy (usukidoll):

w t fu u fuu

OpenStudy (usukidoll):

well if everyone scores bad, teacher's fault ^^

OpenStudy (anonymous):

im not here to argue

OpenStudy (usukidoll):

o-0 no one is arguing

OpenStudy (anonymous):

@anair , please follow me and I will prove it for you.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

yes thats correct.

OpenStudy (anonymous):

v= c1 u1 + c2 u2 + c3 u3, suppose a v + b u1 + d u2=0 then a (c1 u1 + c2 u2 + c3 u3) +b u2 + d u2=0 (a c1 +b)u1 + (a c2+d )u2 + a c3 u3=0\\ a c1 + b =0 a c2 + d =0 a c3=0, since c3 is not zero, then a=0 a= 0 implies 0 c1 + b=0, then b =0 0 c2 + d=0 then d=0 We are done. Actually no need to consider case1 or case 2

OpenStudy (anonymous):

a (c1 u1 + c2 u2 + c3 u3) +b u1 + d u2=0

OpenStudy (anonymous):

There was a misprint in the third line bu2 should be b u1

OpenStudy (anonymous):

Did you understand it?

OpenStudy (anonymous):

yes i understand it much better.

OpenStudy (anonymous):

Excellent

OpenStudy (anonymous):

Sorry, I had to go to dinner, you cold have finished earlier.

OpenStudy (anonymous):

Its ok thank you for ur help!

OpenStudy (anonymous):

YW

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Is contrapositive too confusing?

OpenStudy (anonymous):

yea it was